OpenStudy (anonymous):

Solve 9c + 4 = −23 −3 3 −2.1 4

OpenStudy (diyadiya):

Subtract 4 from both sides What do you get now ?

OpenStudy (diyadiya):

9c +4 -4 = -23 -4

OpenStudy (anonymous):

9c+4=-23 9c=-27 c=-3

OpenStudy (anonymous):

@Diyadiya beat ya!!!!!

OpenStudy (diyadiya):

@ashwinjohn3 Please allow @staceyyyy1 to reply

OpenStudy (mayliz):

9c+4=-23 9c=-23-4 9c=-27 c=-27/9 c=-3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):